# Delayed takedown of illegal content on social media makes moderation ineffective

### v2

@@ Results
Second, we wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries. To estimate a plausible range for $p$, let us analyze the Future of Free Speech policy report (34). The report includes comment data collected from Facebook and YouTube across Germany, France, and Sweden during two weeks in June–July 2023. It also includes the fraction of comments that disappear ($\frac{\text{disappear}}{\text{total}}$) and the proportion of disappeared comments that are legal ($\frac{\text{legal}}{\text{disappear}}$). From this, we estimate the fraction of comments that are illegal: $\frac{\text{illegal}}{\text{total}}=\frac{\text{disappear}}{\text{total}}\times\frac{\text{illegal}}{\text{disappear}}=\frac{\text{disappear}}{\text{total}}\times(1-\frac{\text{legal}}{\text{disappear}})$. Table 1 shows that this apprach yields a range between 0.00002–0.009. We run experiments using these estimates (except for the smallest value, which would require a network with infeasibly large size, as discussed in Methods). In our main analysis we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. 5a for different $p$ values from Table 1. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. 5b the reduction in the prevalence of illegal content for different values of $p$ from Table 1. We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure 5: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations using $\tau=2$. Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1. Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2$ (top) and $\tau=8$ (bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of experiments with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
Last, we examine the effect of network structure by comparing experiments using the empirical network and a synthetic network (see Methods). There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig. 6).
Figure 6: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2$ (left) and $\tau=8$ (right). The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
@@ Discussion
This paper reveals how delays in illegal content takedown impact its effectiveness. Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays. While platforms like TikTok typically respond within a week, illegal content often remains online for over a month on platforms like Facebook, Instagram, Snapchat, and YouTube. Our agent-based model calibrated on this data shows that long delays can substantially diminish the impact of content removal. Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds. Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
@@
These findings have practical implications for social media moderation and regulation. By adopting more proactive takedown policies, platforms may reduce moderation delay and limit the spread of illegal content. Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards. With human flaggers being a limited resource (35), efforts should be directed towards prioritizing the most time-sensitive cases.
@@
For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, the practical impact of such deadlines should be carefully assessed in regulatory debates for two reasons. First, regulatory takedown deadlines can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria. These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged. To better inform policy, platforms should be required to report not only the timestamps for publication and removal of content, but also the detection/notification timestamps to the DSA Transparency Database. This would allow to identify how long content has circulated in the network undetected as well as the moderation delay in response to notifications.
Second, tight takedown deadlines may reduce the time necessary for properly assessing the (un)lawfulness of content, increasing the risk of wrong moderation decisions and thus lowering the quality of content moderation. Flagged content requires careful verification, including report reviews and appeal resolutions, before removal decisions can be made. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects. For example, in France, the Loi Avia was struck down by the Constitutional Council (36) due to concerns about restrictions on freedom of expression (22, 23). These considerations suggest that platforms should remove immediately only obviously illegal content. For cases requiring further assessment, platforms should take time for thorough review. Until the assessment is completed, platforms may implement alternative interventions to reduce the circulation of potentially illegal content.
@@
While our findings are robust, several limitations warrant consideration. The lack of empirical data on illegal content results in several assumptions about the content production process in our model. First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material. We assume static probability distributions of sharing such content, though in reality, this behavior may evolve in response to moderation policies, potentially declining under stricter or faster moderation. Second, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content. If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa. Lastly, we triangulate data from multiple sources, which may introduce inconsistencies. For instance, user activity is calibrated using global Twitter data, whereas the DSA database focuses on European users. This mismatch could lead to unknown biases in our calibration.
@@
Our model of the moderation process also has limitations. The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process. Our model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types. We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries. Our model does not consider such a differentiation.
@@
The database itself has limitations. It relies heavily on voluntary reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms (31). This highlights a broader issue with self-reporting systems. Moderation decisions are shaped by platform-specific policies and guidelines, which may not always align with legal standards: content labeled as “illegal” may in fact be lawful but removed due to internal platform rules. The issue of false positives may have inflated the inferred probability of illegal content from the DSA-TDB.
@@
Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions (37, 38). One could also extend the model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.

### v3

@@ Methods
Modeling social media diffusion
We use SimSoM Truong et al. (2024), an agent-based model that mimics information diffusion on social networks such as Twitter/X, Mastodon, Bluesky, Threads, or Instagram. The platform is represented as a directed network, where nodes correspond to user accounts and edges denote follower relationships, which need not be reciprocal. As in real-world platforms, information circulates through messages that appear on user news feeds. At each time step, a user can post new original messages, reshare content from their feed, or take no action. Messages represent information that could take the form of text, links, hashtags, images, or other media.
To select a message for resharing, an agent samples from the news feed, which contains distinct messages shared by their friends. Once shared by an agent, a message appears in the news feeds of their followers, allowing for propagation across the network (see Discussion). Thus, the exposure of the message depends on its author’s follower count — agents with more followers will achieve greater exposure. The model captures mechanisms of real-world ranking algorithms and sharing behavior, according to which agents prefer to reshare messages that are appealing, recent, and popular Twitter (2023); Truong et al. (2024).
Building on the base model, we incorporate additional features to examine the diffusion of illegal content. Specifically, each message includes a binary attribute indicating whether the content is illegal. Different types of illegal content may be more or less appealing. For example, CSAM is criminally and morally condemned, therefore resharing is expected to be rare. More common forms of illegal content, such as copyrighted media, tend to be highly engaging. Lacking data on the correlation between legality and appeal, we make the simplifying assumption that the two are independent.
We also extend SimSoM to model the heterogeneous activity of social media users, where most agents have low activity levels, and only a few are hyperactive. We model this by assigning an activity level $a$ to each user at the beginning of the simulation. We draw this activity from a power law $P(a)\sim a^{\gamma}$. We estimate the exponent $\gamma=2.85$ by fitting empirical measurements of average daily Twitter activity, based on 81,633,118 posts by 37,882,999 accounts collected over nine days in January 2023. If $a<1$, the user (re)shares a single message with probability $a$ at each time step; else the user (re)shares $\lfloor a\rfloor$ messages.
Modeling illegal content diffusion and removal
To understand the dynamics of illegal content removal, let us consider the lifespan of such content on a platform. While some illegal content is immediately detected and removed using automated detection and decision systems, other is missed due to detection limitations. In such cases, content moderation is delayed until the platform either detects the content itself or receives an external notification. Additional delays may occur when platforms assess the content against their community standards and national laws. If a violation is confirmed, either the content is removed globally or access is restricted in certain jurisdictions.
While real-world moderation involves multiple detection and decision stages, our objective is to capture the aggregate lifetime during which illegal content can diffuse through the network. To this end, we model removal using a stochastic survival process. We make the simplifying assumption that at any time step, illegal content survives with probability $p_{s}$ and is removed with probability $1-p_{s}$. Under this assumption, the probability that content survives for at least $t$ time steps is $P(t)\approx p_{s}^{t}$, since the content must survive $t$ consecutive removal opportunities.
Empirically, takedown delays are well described by an exponential survival curve, $P(t)\approx e^{-t/\tau}$, where $\tau$ is the expected takedown delay or mean lifetime obtained by fitting the DSA-TDB data (see Long takedown delays across platforms). To ensure that the simulation mirrors the same survival curve and $\tau$ values as observed in the empirical data, we parameterize the discrete-time survival process to match this exponential form at integer time steps. Since $P(t)\approx p_{s}^{t}$ in the simulation, we set $p_{s}=e^{-1/\tau}$. Thus, at each time step we remove content with probability $1-e^{-1/\tau}$. When a message is removed, it is deleted from all user feeds.
Simulation framework and metrics
Having specified the diffusion and removal dynamics of illegal content, we now turn to the simulation framework used to evaluate the effects of removal delays — the network structure, outcome metrics, convergence criterion, and parameter settings used in the simulations.
Simulations are conducted on a follower network derived from empirical Twitter data. The network connects users obtained from a 10% random sample of public tweets Nikolov et al. (2020). The dataset comprises users who posted links to news articles, with likely automated accounts excluded. The original follower network constructed from this dataset has 58,29658{,}296 nodes and 10,632,02310{,}632{,}023 edges. We reduce the size of this network to speed up the simulations reported in Takedown fails to curb illegal content prevalence and exposure. We employ $k$-core decomposition to identify N=10,006N=10{,}006 nodes that constitute the $k=94$ core. Subsequently, a random selection of edges is removed to match the density of the original network. This procedure produces E=1,809,798E=1{,}809{,}798 edges, with each node having an average of roughly 180 friends or followers. The main results are robust with respect to the structure of the network (see Appendix, Robustness).
We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, which can be quantified by multiple metrics. Illegal content prevalence is defined as the average ratio of illegal content in circulation at time $t$: $I_{t}=\frac{1}{N}\sum_{i=1}^{N}f_{i,t}$ where $f_{i,t}$ is the fraction of illegal content in user $i$’s feed at time $t$ and $N$ is the number of agents. The more illegal content is present in the system, the more harm it can potentially cause. However, a piece of illegal content may be removed from a user’s feed before the user sees it. Therefore we also measure potential impact more directly using two exposure metrics. Impressions measures the number of times illegal content is seen by users. Note that a single user can be exposed to many posts with illegal content, or even the same post multiple times. Reach is a complementary measure of exposure, capturing the number of unique users who see any amount of illegal content. For each of these metrics, we calculate the reduction percentage relative to a baseline measure without any content removal. For example, the reduction in illegal content prevalence is calculated by $(\langle\bar{I}_{\text{baseline}}\rangle-\bar{I}_{\text{removal}})/\langle\bar{I}_{\text{baseline}}\rangle$, where $\bar{I}$ is the illegal content prevalence at the end of a simulation and $\langle\bar{I}_{\text{baseline}}\rangle$ is the baseline average.
For each set of parameters, we run 70 simulations without any content removal; these are averaged to obtain baseline metrics. We then run 70 simulations with content removal; illegal content prevalence and exposure metrics (reach and impressions, defined in Takedown fails to curb illegal content prevalence and exposure) are averaged across these runs. All simulations start from random conditions.
Results
Long takedown delays across platforms
Figure 2: Complementary cumulative distribution functions (CCDFs) of time delay in illegal content takedown on five platforms. Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
Our analysis of the DSA-TDB for five major social media platforms reveals substantial variation in illegal content takedown delays, ranging from several days to weeks and in some cases up to several years. Figure 2 presents the empirical distributions of takedown delays derived from DSA-TDB Statements of Reasons for the five platforms. By fitting exponential survival models of the form $P(t)\sim e^{-t/\tau}$ to these empirical curves, we estimate expected takedown delays $\tau\approx 31$ days for Facebook, 87 days for Instagram, 286 days for YouTube, 6 days for TikTok, and 136 days for Snapchat. The exponential models provide a reasonable approximation for short delay periods but diverge from the empirical curves at longer times, where heavier tails are evident. Consequently, these estimates of $\tau$ should be interpreted as optimistic summaries, as the exponential specification downweighs rare but prolonged takedown delays observed in the data. To account for both cross-platform variation and uncertainty in the tail behavior of moderation delays, our simulations (Section Takedown fails to curb illegal content prevalence and exposure) explores a broad range of $\tau$ values that both encompasses and extends beyond the empirically observed estimates, up to a maximum of $\tau=739$ days.
Table 2: Expected content takedown delay $\tau$ of illegal content by type and platform.
YouTube TikTok
Illegal Content Type Observations Delay Observations Delay
Copyright violation 375,531 276
Privacy violation 17,737 2,362 14,174 8
Other illegal 9,076 58 15,375 7
Counterfeit 2,500 78
Circumvention 1,174 769
Hate speech 763 12 19,036 10
Defamation 718 36 1,225 5
Child endangerment 647 88 11,881 1
Court order 254 174
Non-consensual intimate imgs 12,271 6
Fraud 4,922 54
Harassment 4,853 4
Terrorism 4,158 8
National security 1,265 47
CSAM 1,249 5
Variation in takedown delays within platforms may further reflect differences across types of illegal content, as some violations can be assessed quickly while others require more extensive investigation. Table 2 reports estimated $\tau$ values by illegal content type for YouTube and TikTok. Corresponding empirical survival curves for each content type are shown in Appendix, Empirical expected takedown delay by illegal content type. The remaining three platforms are excluded from this analysis because their SoRs do not provide sufficiently informative classifications of illegal content types. Both YouTube and TikTok exhibit substantial heterogeneity across content categories: on YouTube, estimated delays range from 12 to 2,362 days, whereas on TikTok they range from one to 54 days. While this variation underscores that illegal content moderation is far from uniform, we next simulate a range of takedown scenarios designed to encompass most of the empirically observed values of $\tau$.
Takedown fails to curb illegal content prevalence and exposure
Based on the above empirical analysis, we simulate various takedown scenarios by exploring a broad range of expected content takedown delay values, 2 hours $\leq\tau\leq 739$ days. We measure the effects of time delay in content takedown by the reduction of illegal content throughout the network, as quantified by the prevalence and exposure metrics defined in Methods.
Figure 3: Impact of takedown delay on illegal content. The reduction in the prevalence (pink), reach (green), and impressions (purple) of illegal content is plotted as a function of the expected takedown delay $\tau$. Shading represents the 95% confidence intervals estimated using non-parametric bootstrapping. The vertical lines indicate the expected takedown delays obtained by fitting the DSA-TDB data for different platforms.
Fig. 3 plots the reduction in illegal content prevalence, impressions, and reach as a function of the expected takedown delay. Content removal has maximal impact under rapid takedowns. For expected delay below nine hours, removal results in approximately 95–100% reduction of illegal content across all measures. Moderation becomes less effective as the expected time delay increases, allowing illegal content to spread before being removed. Between approximately 9 hours–46 days, the reduction in prevalence of illegal content drops from 95% to negligible levels. The empirical values measured from DSA-TDB data for TikTok and Facebook are within this interval. Those for Instagram, Snapchat, and YouTube correspond to longer delays, where removal has no significant effect on illegal content prevalence. Although the overall trend is similar across the two exposure metrics, the reduction in impressions and reach drops even faster for delays of up to a few days. For example, when illegal content is removed within a day on average, reach and impressions are only reduced by about 60%. In summary, these findings show that delays beyond a day significantly diminish the effect of content removal, and that all platforms have longer delays, even in our optimistic estimates.
@@ Discussion
This paper reveals how delays in illegal content takedown affect prevalece and exposure of illegal content. Using empirical takedown data for five social media platforms from the DSA-TDB, we find large variations in average takedown delays. While TikTok typically responds within a week, illegal content often remains online for over a month on Facebook, Instagram, Snapchat, and YouTube. Our agent-based model calibrated on this data shows that such long delays can substantially diminish the impact of content removal. Although this relationship may appear intuitive, our analysis reveals non-linear dynamics between takedown speed and moderation effectiveness, including critical delay thresholds. Specifically, we show that removing illegal content within one day significantly reduces its prevalence and exposure, whereas prolonged delays allow illegal content to permeate user feeds — beyond 46 days, takedown becomes futile in curbing the spread of illegal content.
These findings have practical implications for legislation and social media moderation. For legislation, the importance of timely moderation appears to support the argument for implementing tight illegal content removal deadlines, as imposed under NetzDG in Germany and KoPl-G in Austria prior to the DSA. However, tight takedown deadlines may have detrimental effects on moderation quality and freedom of speech. Takedown deadlines reduce the time necessary to properly assess the (un)lawfulness of content, which often requires careful verification, including report reviews and appeal resolutions. Furthermore, definitions of what content is illegal differ across countries, demanding careful assessment of national context to avoid over-removal. On the other hand, legislation that defines tight takedown deadlines may impose fines for platforms that systematically fail to meet them. The threat of such fines may incentivize platforms to over-remove content to avoid sanctions, thereby stifling freedom of speech online Europe and others (2017). In France, the Loi Avia was struck down by the Constitutional Council Conseil Constitutionnel (2020) precisely because of concerns about restrictions on freedom of expression Vie publique (2020); Heldt (2020). The Constitutional Council argued that the determination of the unlawfulness of the content is often not based on its obvious nature but is a matter for assessment by competent authorities. The stringent takedown deadline granted to the platforms for enforcement did not allow operators to wait for a decision by a judge, thus violating freedom of expression. Any regulatory framework must ensure effective moderation of illegal content while remaining practical for platforms, and mindful of potential adverse effects.
@@
These considerations temper the implications of our results for proactive moderation policies. Even though reducing takedown delays could significantly limit the spread of illegal content, platforms should remove immediately only obvious cases. For cases requiring further assessment (e.g., borderline content, national legislation requirements), platforms should take time for thorough review to avoid hasty decisions and over-removal. Until the assessment is completed, platforms may implement alternative interventions, such as downranking or reduced visibility, to reduce the circulation of potentially illegal content without resorting to premature removal. Another approach to improving content moderation is to strengthen collaboration with trusted flaggers — individuals or organizations that identify and report content violating community guidelines or legal standards. With human flaggers being a limited resource Facebook (2024), efforts should be directed towards prioritizing the most time-sensitive cases.
@@
Another issue arising from strict regulatory takedown deadlines is that they can only address the timespan between the notification of illegal content and its removal, not the period during which illegal content circulates in the network undetected. Therefore, the impact of tight deadlines on the spread of illegal content is inherently limited. In fact, most platforms already report fast reactions after notifications: a median time below 24 hours according to 2024 DSA Transparency Reports and removal of 80% of illegal content within 24 hours according to 2022 national Transparency Reports for Germany and Austria. These figures suggest that a large portion of the long delays measured from the DSA-TDB data occurs before illegal content is detected or flagged.
To further increase the utility of the DSA Transparency Database for research and policy on content moderation, platforms should be required to report the timestamps not only for publication and removal of content, but also for detection/notification. This would make it possible to identify how long content has circulated in the network undetected, and to evaluate if platforms do react timely to violations and notifications. Moreover, SoRs for content and account restrictions should include additional indicators of reach and scale. Simple engagement metrics (number of followers, views, reshares) would enhance interpretability and support monitoring the spread of illegal content. Finally, individual SoRs should be linked to accounts in a privacy-preserving way to enable identification of the volume of repeat offenders Groesch et al. (2026).
@@
While our findings are robust, several limitations warrant consideration. The lack of empirical data on illegal content results in several assumptions about the content production process in our model. First, the model focuses on the spread of content already posted, without accounting for the user incentives or legal risks influencing decisions to post illegal material. Second, we assume static probability distributions of sharing such content. Future research could extend our model by accounting for the fact that users often alter their posting patterns in response to moderation decisions, potentially declining under stricter or faster moderation Srinivasan et al. (2019); Jhaver et al. (2019). Third, as mentioned in Methods, we assume uniform appeal and spreading potential across all types of content. If illegal content is generally more engaging, the model may underestimate the impact of delayed takedown, and vice versa. Fourth, we triangulate data from multiple sources, which may introduce inconsistencies. For instance, user activity is calibrated using global Twitter data (in particular tweets linking to news), whereas the DSA-TDB focuses on European users. This mismatch could lead to unknown biases in our calibration. Lastly, our extension of SimSoMneglects out-of-network content that could be recommended in user feeds. Further experiments could explore the effects of out-of-network recommendations on illegal content prevalence and exposure.
@@
Our model of the moderation process also has limitations. The DSA-TDB data only captures the downstream outcomes of content moderation, neglecting the detection process. The model thus treats all illegal content as equally detectable, which may overlook important variation in the detection of different content types. We also make the simplifying assumption that when illegal content is removed, it is deleted from all user feeds. In practice, however, some content may be illegal only in certain jurisdictions and access may be selectively disabled for users in these countries. We do not consider such a differentiation.
@@
The database itself has limitations Trujillo et al. (2025); Groesch et al. (2026). It relies heavily on self-reporting by platforms, leading to inconsistent coverage and reporting quality both across and within platforms. This highlights a broader issue with self-reporting systems. Moderation decisions are shaped by platform-specific moderation guidelines and practices. Classification of content as illegal does not follow coherent and reliable procedures across all platforms. Platforms are not required to assess lawfulness of content. Moderation is typically guided by a platform’s terms and conditions to avoid the cost and complexity of legal review Keller (2023). As a result, content classified as illegal may neither reflect actual unlawfulness nor capture all illegal content Groesch et al. (2026). These issues may have affected the illegal content probability inferred from the DSA-TDB.
@@
Future research could extend our model to compare the effects of takedown versus alternative types of interventions, as well as explore combinations of different intervention instruments. Finally, future models should consider the trade-offs between the harm caused by persistent illegal content and the unitended effects of moderation tools, e.g., the limitation of free speech caused by false-positive errors.
@@ Acknowledgements
Appendix
Alternative platform compliance data
The Transparency Reports that platforms are required to publish under the DSA (Art. 15) are a potential starting point to understand how content takedown is practiced across different platforms. They include metrics such as “the median time needed for taking the action.”
However, despite formal compliance, reporting practices differ significantly across platforms, preventing thorough comparative analysis. Moreover, median turnaround times in transparency reports only cover the time between notification and removal of content. As a result, they omit the substantial period during which problematic content circulates in the network prior to notification. Finally, these reports do not provide moderation delays specific to illegal content, and platforms do not allow external access to the underlying data, preventing independent validation.
To overcome these limitations, we instead rely on the DSA Transparency Database (DSA-TDB), as described in the main text.
Illegal content probability modeling
As there is limited access to data on illegal content, it is difficult to determine how such individual probability of posting illegal content is distributed across users. We consider three alternative distributions of individual-level illegal posting probabilities that all share the same population mean $p=0.01$.
First, we consider a nearly homogeneous scenario in which all users have similar propensities to post illegal content. In this case, $p$ is drawn from a normal distribution with small variance, $P(i)\sim\mathcal{N}(\mu=0.01,\sigma=0.001)$. Second, we consider a skewed but unimodal distribution that allows for moderate heterogeneity. We model this using a beta distribution, $P(i)\sim i^{\alpha-1}(1-i)^{\beta-1}$, where the mean is given by $\frac{\alpha}{\alpha+\beta}$. We set $\alpha=10$ and $\beta=990$, yielding the same mean probability $p=0.01$.
Third, we explicitly model population heterogeneity by distinguishing between high-risk and low-risk accounts. This reflects the intuition that a small fraction of users may repeatedly post severe illegal content (e.g., CSAM, terrorist propaganda, or scams), while the majority of users only occasionally violate the law, for example by sharing copyrighted material. In this two-group model, a fraction $s_{H}$ of users belong to the high-risk group, while the remaining $1-s_{H}$ belong to the low-risk group. Within each group, individual probabilities are again drawn from beta distributions as described above. The average illegal content probability across the full population is thus $p=s_{H}\frac{\alpha_{H}}{\alpha_{H}+\beta_{H}}+(1-s_{H})\frac{\alpha_{L}}{\alpha_{L}+\beta_{L}}$, where the $\alpha$ and $\beta$ subscripts indicate the groups. We set $s_{H}=0.1$ to reflect a minority of high-risk accounts. We use the values of $\alpha_{L}=0.1$, $\beta_{L}=90$ for the low-risk group and $\alpha_{H}=3$, $\beta_{H}=30$ for the high-risk group, resulting in $p=0.01$.
We explore whether the main outcome variable is sensitive to these different assumptions about illegal content distributions. The reduction in illegal content prevalence shows no statistically significant differences, suggesting that the choice of underlying distribution does not affect the main findings (see Fig. S1 in Robustness). The simulations discussed in the main text use the two beta distributions. We additionally show that our findings are robust to the relative sizes of the groups (see Fig. S2 in Robustness).
Figure S1: Robustness of results with respect to different distributions of $p$ for $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom) based on 10 simulations. The illegal content reduction does not vary significantly ($P\geq 0.13$ across pairs of simulations with different distributions, using Mann–Whitney U tests with Bonferroni correction).
Robustness
The results presented in the main text may depend on the distributions of individual-level illegal posting probabilities, the proportion of high-risk accounts $s_{H}$, the illegal content probability $p$, and the underlying follower network. We therefore investigate if the results remain valid when changing these parameters. For these robustness checks, we use two different values of expected content takedown delay, $\tau=2.89$ and $\tau=11.54$, corresponding to median delays of 2 and 8 days, respectively, and assign users to the high- and low-risk groups randomly in each run.
Fig. S1 shows that the main results are robust with respect to different assumptions about the distributions of individual illegal posting probabilities.
For the following analyses varying high-risk group size and illegal content probability, we need a large follower network to generate stable interaction dynamics using low values of these parameters ($s_{H}=10^{-3}$ and $p=3\times 10^{-4}$). Therefore, we use the original network but remove a random subset of edges to speed up the simulations. The average in/out-degree is thus reduced to 18 friends per node, resulting in a network with N=58,296N=58{,}296 nodes and E=1,063,202E=1{,}063{,}202 edges.
Figure S2: Robustness of results with respect to different high-risk group sizes $s_{H}$ for $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom) based on 20 simulations. The illegal content reduction does not vary significantly ($P\geq 0.72$ across pairs of simulations with different $s_{H}$ values, using Mann–Whitney U tests with Bonferroni correction).
Let us examine the sensitivity of the results to the size of the high-risk group. While our baseline setting assumes $s_{H}=0.1$, we explore several orders of magnitude smaller values of the high-risk group size ($10^{-3}\leq s_{H}\leq 10^{-1}$) to capture scenarios in which illegal content production is concentrated among a very small fraction of users. The main outcome variable, reduction in illegal content prevalence, shows no statistically significant differences (Fig. S2), suggesting that our findings are robust to changes in $s_{H}$.
We wish to explore how the main result is affected by the volume of illegal content, which varies across platforms and countries. We run simulations with the empirical estimates of $p$ reported in Table 1 (except for the smallest value, which would require a network with infeasibly large size). In the main text, we report the reduction in illegal content prevalence, $1-\frac{\bar{I}_{\text{removal}}}{\langle\bar{I}_{\text{baseline}}\rangle}$. To show that this ratio is independent of $p$, we plot the scaling relationship between $\bar{I}_{\text{removal}}$ and $\bar{I}_{\text{baseline}}$ in Fig. S3a for different $p$ values from Table 1. We observe only a weak sublinear relationship, $\bar{I}_{\text{removal}}\sim\bar{I}_{\text{baseline}}^{0.94}$, suggesting that the ratio is nearly independent of $p$. We also plot in Fig. S3b the reduction in the prevalence of illegal content for different values of $p$ from Table 1. We observe more noise for small $p$, but none of the differences between the values in our main results ($p=0.01$) and those for smaller $p$ values are statistically significant. These analyses suggest that our main results are robust with respect to the volume of illegal content in the system.
Figure S3: Robustness analysis for illegal content probability. (a) Scaling between illegal content prevalence in removal and baseline conditions. Each data point represents the proportion of illegal content in the system with and without removal. Results are based on 40 simulations using $\tau=2.89$ (median takedown delay of two days). Each simulation is done with an empirical estimate of $p$ (vertical lines) for Facebook and YouTube users from Germany, France, and Sweden, derived from the data reported in Table 1. Estimates of $p$ are based on the proportion of disappeared content that is subsequently judged to be legal. The gray line represents a least-squares fit of the log-transformed data, yielding a slope of $0.94\pm 0.01$. (b) Illegal content reduction for different illegal content probabilities corresponding to the same countries and platforms, based on 10 simulations with $\tau=2.89$ (median takedown delay of two days, top) and $\tau=11.54$ (median takedown delay of eight days, bottom). The result does not vary significantly ($P\geq 0.13$ across pairs of simulations with different $p$ values, using Mann–Whitney U tests with Bonferroni correction).
Last, we examine the effect of network structure by comparing simulations using the empirical network and a synthetic network. We create a synthetic network with features ubiquitous in real-world networks, namely, the presence of hubs and clustering (directed triads). To this end, we use a directed variant of the random-walk growth model Vázquez (2003). We configure the parameters of this model to generate a network comparable in size and average degree to the network used in the main experiment. In particular, we initialize the network with 181 fully connected nodes and assume new nodes have fixed out-degree $k_{out}=180$. New nodes are added by first randomly selecting one node to follow. For the remaining $k_{out}-1$ connections, each new node has an equal probability to follow a random friend of the initially chosen node or a randomly selected node in the network. This results in a network with N=10,006N=10{,}006 nodes and E=1,810,905E=1{,}810{,}905 edges. There is no statistically significant difference in the reduction in illegal content prevalence between these two networks, indicating the robustness of our findings (Fig. S4).
Figure S4: Robustness analysis with respect to network structures based on 10 simulations with $\tau=2.89$ (median takedown delay of two days, left) and $\tau=11.54$ (median takedown delay of eight days, right). The illegal content reduction does not vary significantly ($P=0.11$ and $P=0.22$, respectively, using Mann–Whitney U tests with Bonferroni correction).
Empirical expected takedown delay by illegal content type
Figure S5: Distributions (CCDF) of time delay in takedown on YouTube by type of illegal content. Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
Figure S6: Distributions (CCDF) of time delay in takedown on TikTok by type of illegal content. Each plot reports empirical data from the DSA-TDB (green dots) along with an exponential fit (purple line).
We wish to examine the variation in takedown delays across types of illegal content (Table 2). Fig. S5 and Fig. S6 show empirical survival time curves for different illegal content types from DSA-TDB SoR records of YouTube and TikTok. By fitting exponential survival time curves $P(t)\sim e^{-t/\tau}$ to this empirical data, we estimate the ranges of $\tau$ for these platforms: 12 days $\leq\tau\leq$ 2,362 days for YouTube and 1 day $\leq\tau\leq$ 54 days for TikTok.